# On structural contraction of biological interaction networks

### v3

@@ 4.2.2 General results on strict contraction
We state our result next which is can be proved in a similar manner to the argument in Example 2 in §4.2. All the testes stipulated in the statement of the theorem can be carried out computationally.
@@
Let $K\subset\mathscr{C}_{\bar{x}}\cap\mathbb{R}_{+}^{n}$ be any compact set, and assume that $\varphi(t,K)$ is uniformly bounded (which follows automaically from the existence of a steady state or a conservation law as shown in Corollary 3.13). Then for any two solutions $x_{1}(\cdot)$ and $x_{2}(\cdot)$ of $\dot{x}=\Gamma R(x)$ that start in $K$, there exists $\bar{\theta}>0$, $c>0$ such that
$|x_{1}(t)-x_{2}(t)|_{PB}\;\leq\;e^{-ct}|x_{1}(0)-x_{2}(0)|_{PB}\quad\mbox{for all}\;t\geq 0,$
where $P$ is the corresponding contractor matrix defined before Lemma 4.6.
@@
Let $K\subset\mathscr{C}_{\bar{x}}\cap\mathbb{R}_{+}^{n}$ be any compact set, and assume that $\varphi(t,K)$ is uniformly bounded. Then for any two solutions $x_{1}(\cdot)$ and $x_{2}(\cdot)$ of $\dot{x}=\Gamma R(x)$ that start in $K$, there exists $c>0$ such that
@@ 5.1 Preliminaries
Consider a dynamical system $\dot{x}=f(t,x)$, where $f$ is defined on an convex forward invariant set $K\subset R^{n}$, is differentiable in $x$, and is continuous in $t$. Let |.||.|| . | be a norm on $\mathbb{R}^{n}$. Then, $f$ is said to be contractive after a small overshoot and short transient (CSOST) on $K$ w.r.t to |.||.|| . | if for each $\varepsilon>0$ and each $\tau>0$ there exists $c=c(\tau,\varepsilon)>$ such that
$|\varphi(t+\tau;t_{0},a)-\varphi(t+\tau;t_{0},a)|\leq(1+\varepsilon)e^{-(t-t_{0})c}|a-b|,$
@@ 5.2 Contraction after small transients and overshoots
    $R(x,t)$ is continuousely differentiable in $x$, and continuous in $t$.
@@
For uniformly bounded trajectories, persistence results generalize to the time-varying system (29) when the network admits no critical siphons [14]. In our context, we need to establish a uniform bound
@@
Let ${\mathscr{N}}$ be a given network, and it assume that it admits trivial siphons only. Consider a fixed choice of admissible kinetics $R$, and consider the corresponding dynamical system (29). Fix a proper stoichiometric class $\mathscr{C}$ and let $K\subseteq\mathscr{C}$ be compact and forward invariant. Then, for all $\tau>0$, $\varphi(\tau,K)$ is a compact subset of the open positive orthant.
@@
Fix $\tau>0$. Since $\varphi$ is continuous, then $\varphi(\tau;K)$ is compact. To prove that it is a subset of the open positive orthant, assume by contradiction that $\exists x_{1}$ such that $x_{1}\in\varphi(\tau;K)\cap\partial\mathbb{R}_{+}^{n}$. Let $x_{0}:=\varphi(-\tau;x_{1})$. By the forward invariance of $\mathscr{C}$, we have $x_{0}\in\mathscr{C}$. Note also that $x_{1}=\varphi(\tau;x_{0})\in\partial\mathbb{R}_{+}^{n}$. Next, we claim that there is $\tau_{1}>0$ such that $\varphi(t;x_{0})\in\mathbb{R}_{+}^{n}$ for all $t\in(0,\tau_{1})$. The latter claim can be proven as follows: If no such $\tau_{1}$ exists, then there an interval $J$ such that $\varphi(t;x_{0})\in\partial\mathbb{R}_{+}^{n}$ for all $t\in J$. Hence, there is a face of the boundary that intersects the stoichiometry class for which solutions evolve in such intersection for all $t\in J$. Hence, let us call this face $L_{\Sigma}=\{x:x_{i}=0,i\in\Sigma\}$ for some $\Sigma\subset{\mathscr{S}}$. This means $\dot{x}_{i}=0$ for all $t\in J,i\in\Sigma$, and $\dot{x}_{i}>0$ for all $t\in J,i\not\in\Sigma$. Moreover outgoing reactions in which $i$ is a reactant are such that $R_{j}(x(t),t)=0$. Hence, all reactions that have $i$ as a product need to also have 0 rates (otherwise $\dot{x}_{i}(t)>0$). This implies that all reactions feeding into $i$ have zero rates, hence they have a reactant in $\Sigma$, therefore $\Sigma$ is non-trivial siphon which is a contradiction.
Now, since $\varphi(\tau_{1}/2;x_{0})\in\mathbb{R}_{+}^{n}$, then $\varphi(\tau-\tau_{1}/2;\varphi(\tau_{1}/2;x_{0}))=x_{1}\in\partial\mathbb{R}_{+}^{n}$. This implies that a trajectory starting in the interior reaches the boundary in finite time which contradicts the forward invariance of the open positive orthant.
@@
Let a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$ be given and assume that it admits a GLF ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$, it has only trivial siphons, and is weakly contractive. Let $\mathscr{C}$ be a stoichiometric class. Let $K\subset\mathscr{C}$ be compact and forward invariant. Then, the dynamical system (29) defined on $K$ is CSOST with respect to |.$|.|_{B}$ defined on $\mathscr{C}$.
@@
where ()($\star$)( ⋆ ) follows from nonexpansiveness.
Pick an arbitrary $\varepsilon>0$. If $\varepsilon<\tilde{\varepsilon}_{\bar{\theta}}$, then there exists $\theta\in(0,\bar{\theta})$ such that $\varepsilon=\tilde{\varepsilon}(\theta)$. Hence, using (31) we get:
@@ 6 Biochemical Examples
We represent BINs graphically as Petri-nets [68], where a reaction is a represented by a square, and a species is represnted by a circle. We review next some biochemical motifs which are amenable to our results. Other examples and more complex cascades can be analyzed too as discussed in [5, 1]. All the examples presented in this section are conservative and have only trivial siphons.
[…]
@@ 6.1 Post-translational modification (PTM)
A simplified network was a already presented in (9). Here, we consider the full PTM given below:
@@
Using Corollary 2.15, we find matrices $\Lambda_{1},..,\Lambda_{8}$ via the subroutine FindLambda.m as follows:
@@
Note that $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})=0$. Using the notation defined in §4.2.1, we can see that the network is weakly contractive with $S_{0}=\{1,6\}$. Therefore the contractor matrix is given as $P_{\theta}=\mathop{\mbox{diag}}([1,1+\theta,1+\theta,1+\theta,1+\theta,1])$. The scaled norm is given as $P_{\theta}B$. Remember that $B$ is not unqiue, one such choice provided by the subroutine FindBfromC.m in the package LEARN is:
@@
Hence, for an aribtrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the $P_{\theta}B$-weighted $\infty$-norm, where $\theta>0$ is small enough depending on the choice of the compact set. In addition,the trajectories entrain to periodic time-varying parameters.
@@ 6.2 Three-Body Binding
Consider the three-body network shown in (26) and depicted in Figure 3(b). The reactions are listed in (26), let $x_{1}=[{\rm A}],x_{2}=[{\rm B}],x_{3}=[{\rm C}],x_{4}=[{\rm AB}],x_{5}=[{\rm BC}],x_{6}=[{\rm ABC}].$Using FindLambda subroutine, the matrices $\Lambda_{1},..,\Lambda_{12}$ are given as
@@
Therefore the expression $\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell}$ can be written as follows:
@@
which is strictly negative over any positive compact set. Note $P_{\theta}=I,B=I$. Hence, for an aribtrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the standard $\infty$-norm. In addition, the trajectories entrain to periodic time-varying parameters.
6.3 Kinetic prooreading in $T$-cell activation
The kinetic proofreading model introduced by [46]. It was repurposed by McKeithan [64, 75] to model $T$-cell activation, and is supported experimentally [41]. In this model, a ligand $M$ binds to a $T$-cell receptor. The receptor-ligand complex then undergoes a series of reactions to produce an active complex $C_{N}$. Figure 3(c) depicts the reaction network. The reactions can be written as follows:
$M+L$ $\rightleftharpoons C_{0}\mathop{\longrightarrow}\limits C_{1}\mathop{\longrightarrow}\limits...\mathop{\longrightarrow}\limits C_{N}$
$C_{1}$ $\mathop{\longrightarrow}\limits M+L,C_{2}\mathop{\longrightarrow}\limits M+L,...,C_{N}\mathop{\longrightarrow}\limits M+L.$
@@ 6.3 Kinetic prooreading in $T$-cell activation
The matrices $\Lambda_{1},..,\Lambda_{7}$ can be found, and the expression $\sum_{\ell}\rho_{\ell}\Lambda_{\ell}$ can be written as follows:
@@
Note that $\mu_{\infty}(\sum_{\ell}\rho_{\ell}\Lambda_{\ell})=0$, but the network is weakly contractive with $S_{0}=\{3,5,6\}$ and $S_{-}=\{1,2,4,7\}$. We find a matrix $B$ satisfying $C=B\Gamma$, and the weighting matrix for the norm can be written as:
@@
Hence, for an aribtrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the $P_{\theta}B$-weighted $\infty$-norm, where $\theta>0$ is small enough depending on the choice of the compact set. More exactly, let $\widetilde{K}$ be the minimal compact set of the form (23) that contains $K$. Then, the upper bound $\bar{\theta}$ that appears in Theorem 4.11 can be written as:
@@
$\qquad\qquad\left.\frac{\theta}{\theta+1}(\rho_{6}+\rho_{5}+\rho_{3}+\rho_{2}+\rho_{1}),\rho_{1}+\rho_{2}-\theta(\rho_{3}+\rho_{6}+\rho_{7})\right\}<0,$
@@
Phosphorelays
Unlike advanced organisms which rely on PTMs for signalling, bacteria often use a motif known as phosphotransfer where a histidine kinase provides its phosphate to a response regulator. This motif is a building block in what is known as a two-component signaling systems [79]. Experimental examples include EnvZ/OmpR system in E. Coli [21]. Phosphotransfer cascades are known as phosphorelays [16, 57]. Examples include Bacillus subtilis [45] and yeast [69]. A general network is depicted in Figure 3(d). The reaction network for the case $n=2$ can be written as:
@@
Let $\Lambda_{1},..,\Lambda_{14}$ be the matrices that can found using Corollary 2.15. Let $\Xi=\sum_{\ell}\rho_{\ell}\Lambda_{\ell}$. Then $\Xi$ can be written as:
@@
We have $\mu_{\infty}(\Xi)=0$, but the network can be verified to be weakly contractive with nonexpansivity depth 2. In particular, $S_{-}=\{1,..,8\},S_{01}=\{9,..,14\},S_{02}=\{15\}$. The weighting matrix $P_{\theta}B$ can be written as:
@@
Hence, for an aribtrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the $P_{\theta}B$-weighted $\infty$-norm, where $\theta>0$ is small enough depending on the choice of the compact set. In addition,the trajectories entrain to periodic time-varying parameters.

### v4

@@ 4.2.2 General results on strict contraction
We state our result next which can be proved in a similar manner to the argument in Example 2 in §4.2. All the tests stipulated in the statement of the theorem can be carried out computationally.
@@
Let $K\subset\mathscr{C}_{\bar{x}}\cap\mathbb{R}_{+}^{n}$ be any compact set, and assume that $\varphi(t;K)$ is uniformly bounded (which follows automatically from the existence of a steady state or a conservation law as shown in Corollary 3.13). Then for any two solutions $x_{1}(\cdot)$ and $x_{2}(\cdot)$ of $\dot{x}=\Gamma R(x)$ that start in $K$, there exist $\bar{\theta}>0$, $c>0$ such that
${\forall\theta\in(0,\bar{\theta})},|x_{1}(t)-x_{2}(t)|_{P^{{\theta}}B}\;\leq\;e^{-ct}|x_{1}(0)-x_{2}(0)|_{P^{{\theta}}B}\quad\mbox{for all}\ t\geq 0,$
where $P^{{\theta}}$ is the corresponding contractor matrix defined before Lemma 4.6.
@@
Let $K\subset\mathscr{C}_{\bar{x}}\cap\mathbb{R}_{+}^{n}$ be any compact set, and assume that $\varphi(t;K)$ is uniformly bounded. Then for any two solutions $x_{1}(\cdot)$ and $x_{2}(\cdot)$ of $\dot{x}=\Gamma R(x)$ that start in $K$, there exists $c>0$ such that
@@ 5.1 Preliminaries
Consider a dynamical system $\dot{x}=f({x,t})$, where $f$ is defined on a convex forward invariant set $K\subset{\mathbb{R}}^{n}$, is differentiable in $x$, and is continuous in $t$. Let $|.|$ be a norm on $\mathbb{R}^{n}$. Then, $f$ is said to be contractive after a small overshoot and short transient (CSOST) on $K$ w.r.t. to $|.|$ if for each $\varepsilon>0$ and each $\tau>0$ there exists $c=c(\tau,\varepsilon)>{0}$ such that
$|\varphi(t+\tau;t_{0},a)-\varphi(t+\tau;t_{0},{b})|\leq(1+\varepsilon)e^{-(t-t_{0})c}|a-b|,$
@@ 5.2 Contraction after small transients and overshoots
    $R(x,t)$ is continuously differentiable in $x$, and continuous in $t$.
@@
For uniformly bounded trajectories, persistence results generalize to the time-varying system (29) when the network admits no critical siphons [14]. In our context, we need to establish a uniform bound.
@@
Let $\mathscr{N}$} be a given network, and assume that it admits trivial siphons only. Consider a fixed choice of admissible kinetics $R$, and consider the corresponding dynamical system (29). Fix a proper stoichiometric class $\mathscr{C}$ and let $K\subseteq\mathscr{C}$ be compact and forward invariant. Then, for all $\tau>0$, $\varphi(\tau,K)$ is a compact subset of the open positive orthant.
@@
Fix $\tau>0$. Since $\varphi$ is continuous, then $\varphi(\tau;K)$ is compact. To prove that it is a subset of the open positive orthant, assume by contradiction that $\exists x_{1}$ such that $x_{1}\in\varphi(\tau;K)\cap\partial\mathbb{R}_{+}^{n}$. Let $x_{0}:=\varphi(-\tau;x_{1})$. Note that we have $x_{0}\in\mathscr{C}$ since the system is initialized in $\mathscr{C}$. Note also that $x_{1}=\varphi(\tau;x_{0})\in\partial\mathbb{R}_{+}^{n}$. Next, we claim that there is $\tau_{1}>0$ such that $\varphi(t;x_{0})\in\mathbb{R}_{+}^{n}$ for all $t\in(0,\tau_{1})$. The latter claim can be proven as follows: If no such $\tau_{1}$ exists, then there is an interval $J$ such that $\varphi(t;x_{0})\in\partial\mathbb{R}_{+}^{n}$ for all $t\in J$. Hence, there is a face of the boundary that intersects the stoichiometry class for which solutions evolve in such intersection for all $t\in J$. Hence, let us call this face $L_{\Sigma}=\{x:x_{i}=0,i\in\Sigma\}$ for some $\Sigma\subset{\mathscr{S}}$. This means $\dot{x}_{i}=0$ for all $t\in J,i\in\Sigma$.Moreover outgoing reactions in which $i$ is a reactant are such that $R_{j}(x(t),t)=0$. Hence, all reactions that have $i$ as a product need to also have 0 rates (otherwise $\dot{x}_{i}(t)>0$). This implies that all reactions feeding into $i$ have zero rates, hence they have a reactant in $\Sigma$, therefore $\Sigma$ is a non-trivial siphon which is a contradiction.
Now, since $\varphi(\tau_{1}/2;x_{0})\in\mathbb{R}_{+}^{n}$, then $\varphi(\tau-\tau_{1}/2;\varphi(\tau_{1}/2;x_{0}))=x_{1}\in\partial\mathbb{R}_{+}^{n}$. This implies that a trajectory starting in the interior reaches the boundary in finite time which contradicts the forward invariance of the open positive orthant. $\blacksquare$
@@
Let a network ${\mathscr{N}}=({\mathscr{S}},{\mathscr{R}})$ be given and assume that it admits a GLF ${\widetilde{V}}(r)=\|B\Gamma r\|_{\infty}$, it has only trivial siphons, and is weakly contractive. Let $\mathscr{C}$ be a stoichiometric class. Let $K\subset\mathscr{C}$ be compact and forward invariant. Then, the dynamical system (29) defined on $K$ is CSOST with respect to $|.|_{B}$ defined on $\mathscr{C}$.
@@
where $(\star)$ follows from nonexpansiveness.
Pick an arbitrary $\varepsilon>0$. If $\varepsilon<\tilde{\varepsilon}{({\bar{\theta}})}$, then there exists $\theta\in(0,\bar{\theta})$ such that $\varepsilon=\tilde{\varepsilon}(\theta)$. Hence, using (31) we get:
@@ 6 Biochemical Examples
We review next some biochemical motifs which are amenable to our results. The examples are depicted graphically in Figure 3 using the Petri-net convention [68], where a reaction is represented by a rectangle, and a species is represented by a circle. Other examples and more complex cascades can be analyzed too as discussed in [5, 1]. All the examples presented in this section are conservative and have only trivial siphons.
$R_{2}$$R_{1}$$R_{3}$$R_{4}$$^{+}SS$KinasePhosphatase
@@ 6.1 Post-translational modification (PTM)
A simplified network was already presented in (9). Here, we consider the full PTM given below:
@@
Using Corollary 2.15, we find matrices $\Lambda_{1},\dots,\Lambda_{8}$ via the subroutine FindLambda.m using the command FindLambda($\Gamma$,C):
@@
Note that $\mu_{\infty}(\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell})=0$. Using the notation defined in §4.2.1, we analyze this network manually to find that it is weakly contractive with $S_{0}=\{1,6\}$. Therefore the contractor matrix is given as $P_{\theta}=\mathop{\mbox{diag}}([1,1+\theta,1+\theta,1+\theta,1+\theta,1])$. The scaled norm is given as $P_{\theta}B$. Remember that $B$ is not unique, one such choice provided by the subroutine FindBfromC.m in the package LEARN is:
@@
Hence, for an arbitrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the $P_{\theta}B$-weighted $\infty$-norm, where $\theta>0$ is small enough depending on the choice of the compact set. In addition, the trajectories entrain to periodic time-varying parameters.
@@ 6.2 Three-Body Binding
Consider the three-body network shown in (26) and depicted in Figure 3(b). The reactions are listed in (26), let $x_{1}=[{\rm A}],x_{2}=[{\rm B}],x_{3}=[{\rm C}],x_{4}=[{\rm AB}],x_{5}=[{\rm BC}],x_{6}=[{\rm ABC}].$Using FindLambda subroutine, the matrices $\Lambda_{1},\dots,\Lambda_{12}$ are given using the command FindLambda($\Gamma$,C):
@@
Therefore the expression $\sum_{\ell=1}^{s}\rho_{\ell}\Lambda_{\ell}$ can be written using the command FindLambda($\Gamma$,C):
@@
which is strictly negative over any positive compact set. Note $P_{\theta}=I,B=I$. Hence, for an arbitrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the standard $\infty$-norm. In addition, the trajectories entrain to periodic time-varying parameters.
As a numerical illustration of the results, we depict simulations of the three-body binding system in Figure 4. Figure 4(a) shows the non-expansivity of the norm with $B=I$. With 1000 numerically generated trajectories, the maximum logarithmic norm recorded is $7.1636\times 10^{-5}$. Figure 4(b) shows the trajectories converging to the same steady state in the stoichiometric class. Figure 4(c) shows that all the states entrain to a periodic change in the parameters.
Figure 4: Simulations of the three-body binding network using the norm found by LEARN. (a) A simulation of 1000 trajectories in the stoichiometric class corresponding to $A]_{tot}=[B]_{tot}=[C]_{tot}=5$, for the ODE with mass-action kinetics and all kinetic constants set to 1. (b) The numerically-calculated time-derivative of the $B$-distance between 1000 pairs of randomly generated trajectories for the same ODE system in panel (a). Each pair of trajectories starts in the same stoichiometric class. (c) Time-series plots of all six states with $A]_{tot}=[B]_{tot}=[C]_{tot}=5$ and reaction-rates following mass-action kinetics. The periodically-varying kinetic constants are given as $k_{1}=k_{5}=k_{6}=k_{3}=k_{8}=\sin^{2}(2t),k_{2}=k_{4}=k_{7}=\cos^{2}(2t)$.
6.3 Kinetic proofreading in $T$-cell activation
The kinetic proofreading model introduced was by [46]. It was repurposed by McKeithan [64, 75] to model $T$-cell activation, and is supported experimentally [41]. In this model, a ligand $M$ binds to a $T$-cell receptor. The receptor-ligand complex then undergoes a series of reactions to produce an active complex $C_{N}$. Figure 3(c) depicts the BIN. The reactions can be written as follows:
$M+L$ $\rightleftharpoons C_{0}\mathop{\longrightarrow}\limits C_{1}\mathop{\longrightarrow}\limits\dots\mathop{\longrightarrow}\limits C_{N}$
$C_{1}$ $\mathop{\longrightarrow}\limits M+L,C_{2}\mathop{\longrightarrow}\limits M+L,\dots,C_{N}\mathop{\longrightarrow}\limits M+L.$
@@ 6.3 Kinetic proofreading in $T$-cell activation
The matrices $\Lambda_{1},\dots,\Lambda_{7}$ can be found, and the expression $\sum_{\ell}\rho_{\ell}\Lambda_{\ell}$ can be written using the command FindLambda($\Gamma$,C):
@@
Note that $\mu_{\infty}(\sum_{\ell}\rho_{\ell}\Lambda_{\ell})=0$. We perform weak contractivity analysis manually. The network is weakly contractive with $S_{0}=\{3,5,6\}$ and $S_{-}=\{1,2,4,7\}$. We find a matrix $B$ satisfying $C=B\Gamma$ using FindBfromC.m, and the weighting matrix for the norm can be written as:
@@
Hence, for an arbitrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the $P_{\theta}B$-weighted $\infty$-norm, where $\theta>0$ is small enough depending on the choice of the compact set. More exactly, let $\widetilde{K}$ be the minimal compact set of the form (23) that contains $K$. Then, the upper bound $\bar{\theta}$ that appears in Theorem 4.11 can be written as:
@@
$\hskip 17.00024pt\hskip 17.00024pt\left.\frac{\theta}{\theta+1}(\rho_{6}+\rho_{5}+\rho_{3}+\rho_{2}+\rho_{1}),\rho_{1}+\rho_{2}-\theta(\rho_{3}+\rho_{6}+\rho_{7})\right\}<0,$
@@
6.4 Phosphorelays
Unlike advanced organisms which rely on PTMs for signalling, bacteria often use a motif known as phosphotransfer where a histidine kinase provides its phosphate to a response regulator. This motif is a building block in what is known as a two-component signaling system [79]. Experimental examples include EnvZ/OmpR system in E. Coli [21]. Phosphotransfer cascades are known as phosphorelays [16, 57]. Examples include Bacillus subtilis [45] and yeast [69]. A general network is depicted in Figure 3(d). The BIN for the case $n=2$ can be written as:
@@ 6.4 Phosphorelays
Let $\Lambda_{1},\dots,\Lambda_{14}$ be the matrices that can be found using Corollary 2.15 using the command FindLambda($\Gamma$,C). Let $\Xi=\sum_{\ell}\rho_{\ell}\Lambda_{\ell}$. Then, using the output of the subroutine, $\Xi$ can be written as:
@@
We have $\mu_{\infty}(\Xi)=0$, but the network can be verified manually to be weakly contractive with nonexpansivity depth 2. In particular, $S_{-}=\{1,\dots,8\},S_{01}=\{9,\dots,14\},S_{02}=\{15\}$. The weighting matrix $P_{\theta}B$ can be written as:
@@
Hence, for an arbitrary given positive compact set in a stoichiometric class, the trajectories are strictly contractive with respect to the $P_{\theta}B$-weighted $\infty$-norm, where $\theta>0$ is small enough depending on the choice of the compact set. In addition, the trajectories entrain to periodic time-varying parameters.